OpenStudy (anonymous):

find height of a tree that makes a 40'shadow when5'3" in pole is used to make a shadow of 3'

OpenStudy (anonymous):

x= shadow(s) y=height y1/x1=y/x y/40 = 5'3''/3' y=40' * 5'3''/3' calculate

OpenStudy (anonymous):

\[\frac{63}{36}=\frac{h}{480}\] \[h=\frac{63\times 480}{36}=840\] 840 inch = 70 ft

OpenStudy (anonymous):

easiest i think to work in inches and then convert back.